OpenStudy (anonymous):

use the following rule to find 3^2 + 6^2 + 9^2 + .... + (3n)^2

OpenStudy (anonymous):

\[1^2 + 2^2 + 3^2 + .... + n^2 = 1/6n(n+1)(2n+1)\]

OpenStudy (shubhamsrg):

take 3^2 common out.

OpenStudy (anonymous):

no wait shouldn't i just replace n with 3n on the RHS of the rule??

OpenStudy (shubhamsrg):

That would have worked if you had 1^2 + 2^2 + 3^2 ....... (3n)^2 which is not the case here.

OpenStudy (anonymous):

arrey but the first rule is what i posted, and the second one looks like thi,s if you couldn't make out \[3^2 + 6^2 + 9^2 + .... + (3n)^2\]

OpenStudy (shubhamsrg):

I am not understand your confusion.

OpenStudy (anonymous):

and what is 3^2 but (1*3)^2, and what is 6^2 but (2*3)^2

hartnn (hartnn):

again, you can factor out 3^2 from each term like 9^2 = 3^2.3^2 and you get 3^2(1^2+2^2+...)

hartnn (hartnn):

(2*3)^2 = 2^2*3^2

OpenStudy (anonymous):

yes, i know that :P

hartnn (hartnn):

then your problem is done.

hartnn (hartnn):

=3^2*(1/6)n(n+1)(2n+1) =9*(1/6)n(n+1)(2n+1) =(3/2)n(n+1)(2n+1) got this ?